To 5.0 L ideal gas, 1.0 kJ of heat is added. The gas expands against a constant external pressure of 1.0 bar to a final volume of 8.0 L. What is the change internal energy for the gas (1.0 L.bar = 0.10 kJ) 

A.

0.30 kJ

B.

0.70 kJ

C.

1.3 kJ

D.

1.8 kJ


1
Expert's answer
2021-11-23T00:46:02-0500

8-5 = 3L


= 1× 4.875 × 1.10kj


= 6.9625kJ

6.9625×0.1=

=0.70kJ